random231 (random231):

if f be a decreasing continuous function satisfying f(x+y) = f(x) + f(y) - f(x)f(y) f'(0)=-1

random231 (random231):

then \[\int\limits_{0}^{1}f(x)dx = ?\]

random231 (random231):

@hartnn @ganeshie8

OpenStudy (lastdaywork):

@wio

random231 (random231):

any ideas?

OpenStudy (ikram002p):

hmm im trying but im getting nothing :o

OpenStudy (lastdaywork):

I got \[f(x) = 1 - e^x\]

OpenStudy (ikram002p):

how !

OpenStudy (lastdaywork):

First lemme confirm the answer. @random231 is the final answer (-e) ??

random231 (random231):

no

OpenStudy (lastdaywork):

:(

random231 (random231):

but u did it right ur integration came wrong watch the signs

random231 (random231):

ans is 2-e

OpenStudy (lastdaywork):

Auh..okay..my bad XD

OpenStudy (lastdaywork):

Partially differentiate the given equation w.r.t x f'(x+y) = f'(x) [ 1 - f(y)] put x = 0 and f'(x) = -1 f'(y) = f(y) - 1 Then it is a simple diff equation.. :)

OpenStudy (lastdaywork):

Solution - \[f(x) = Ce^x + 1\] implies f'(0) = C = -1 Hence, \[f(x) = 1 - e^x\]

random231 (random231):

oh ok the tricky part was the partial differentiation!

OpenStudy (lastdaywork):

Yep :)

random231 (random231):

but on wat basis did u hav the idea of using partial diff

OpenStudy (lastdaywork):

& I need to work on my arithmetic.. :(

OpenStudy (lastdaywork):

Experience..I took a drop..so had seen such Q's before ;)

OpenStudy (anonymous):

@random231 They gave you the value of the derivative at \(0\), that that is the basis for finding a derivative.

random231 (random231):

hmm getting it, thanks wio!

OpenStudy (anonymous):

In addition, this property f(x+y) = f(x) + f(y) - f(x)f(y) is inherently to some sort of exponential functions.

OpenStudy (anonymous):

Since \[ e^{x+y}=e^xe^y \]after all.

random231 (random231):

oh dear i didnt notice that thanks wio!!! and lastday!!!

OpenStudy (lastdaywork):

Yea..that makes sense.. I never saw the incentive to use Partial differentiation..just knew I had to use it XD

random231 (random231):

okay